# Exploring Decision-based Black-box Attacks on Face Forgery Detection

### v1

@@ III-B Cross-task Perturbation Initialization
Naturally, exploiting the adversarial transferability of face forgery detectors is a possible strategy. However, in practice, we not only do not have access to the training data of the threat model but also do not know the face forgery generation method it adopts, which greatly limits the possibility of using transfer adversarial perturbations. Since face recognition is a very common face-related task, it may help the attack achieve perturbation initialization. Furthermore, inspired by [41], intermediate features may have high correlations between different face-related tasks. From this perspective, using face recognition models to transfer adversarial perturbations is a feasible initialization strategy.
We assume that if there is a high degree of similarity between the intermediate features of the face recognition model and face forgery models, then attacking the face recognition model is equivalent to attacking the face forgery models. To verify the correlation of intermediate features across face tasks, we conduct an empirical study on FaceForensics++ [18]. We randomly select 256 frames to calculate the intermediate features and then average for calculating cosine similarity. Fig. 2 shows the cosine similarity of intermediate features between face forgery detectors and face cognition models. The middle layer selection of the face recognition model is shown in Table V and the face forgery detectors select the second block unit. According to Fig. 2, we find that face recognition models and face forgery detectors have a high correlation between the intermediate features of the same face because these are all face-related tasks and models capture the general features of faces.
@@
Algorithm 1 illustrates the whole pipeline of the cross-task perturbation initialization, where $\mathrm{sign}$ represents the sign operation on the gradient and $\mathrm{Clip}_{x_{r},\ \xi}(\cdot)$ projects the perturbation $\delta$ to $\xi$-ball in order to satisfy $||\delta||_{\infty}\leq\xi$. The time complexity of CPI is $O(K)$, so the calculation of CPI requires $K$ times of forward and backward propagation of the FR model. In this paper, $K=10$. We test the time to calculate CPI on the A100, which is about $50.34$ms for an image, with almost no additional overhead.
Compared with [41], we leverage the fast sign gradient to generate perturbations efficiently and focus on the high feature correlation between face-related tasks, which is beneficial to the generalizability of our method to adversarial attacks and defenses on more face-related tasks. Face recognition distinguishes the identities of different faces and enables the acquisition of more discriminative features. Therefore, CPI can provide a better initialization for face forgery detection, which ensures a more correct optimization direction and provides the possibility of attacking multiple face-related intelligent systems simultaneously.
III-C Frequency Decision-based Attack
Recent work [36, 16, 17] on face forgery detection has illustrated a significant difference in frequency between real and fake faces. In addition, some work [42, 43, 44] also shows that different models rely on different frequency components to make decisions. Inspired by these works, we consider adding noise directly to the frequency of human faces, aiming to eliminate the difference between real and fake faces in frequency to improve the attack efficiency.
Based on the general framework of decision-based attacks [32, 33], we propose the frequency decision-based attack (FDA) on face forgery detection. Fig. 3 summarizes the pipeline of our proposed attack, including the perturbation initialization, the frequency noise projection, and the random perturbation flip. Specifically, we first generate the cross-task adversarial perturbation $\delta_{init}$ by perturbation initialization and resort to binary search to reduce the perturbation magnitude and guarantee adversarial. In frequency noise projection, we add perturbations in the frequency domain and constrain image quality in the spatial domain. Then, we randomly inverse the sign of the partial perturbation, expecting a better adversarial perturbation. Finally, we iterate in frequency noise projection and random perturbation flip until $c(x_{adv})=0$ and $||\delta||_{\infty}\leq\epsilon$.
@@ III-C Frequency Decision-based Attack
11:  $\delta\leftarrow\mathrm{Clip}_{0,\ u}(\delta_{init})$
@@
Perturbation Initialization. Generally speaking, the perturbation initialization is used in decision-based attacks to make clean examples keep adversarial. Given a face classified as real, we utilize the cross-task perturbation initialization to obtain initial perturbations $\delta_{init}$, which satisfies the adversarial constraint. Note that cross-task adversarial perturbations are imperceptible, so faces after cross-task perturbation initialization remain adversarial. Further, in order to reduce the magnitude of the initial perturbation, we choose to apply the binary search to make the initial adversarial face as close to $x$ as possible. Binary search in perturbation initialization is summarized in Algorithm 2. Given an initial perturbation $\delta_{init}$, we continuously dichotomize the perturbation range from $[0,\max(\delta_{init})]$ on the premise of ensuring adversarial, and finally obtain the final adversarial perturbation $\delta$. In this paper, we set $k$ to 10. With the help of binary search, we obtain the final initial perturbation $\delta$, where $||\delta||_{\infty}\leq\epsilon^{\prime},\epsilon^{\prime}\in(0,\max|\delta_{init}|]$.
Frequency Noise Projection. Here, $\mathcal{D}(\cdot)$ denotes discrete cosine transform (DCT) and $\mathcal{D^{\prime}}(\cdot)$ denotes inverse discrete cosine transform (IDCT). First, we transform the image from the spatial domain to the frequency domain with DCT. Then we add a random noise in the frequency domain $\eta\sim\{-\gamma,\gamma\}^{d}$. Finally, we project it to the $\epsilon^{\prime}$-ball in the spatial domain and gradually decrease $\epsilon^{\prime}$ until it is satisfied $\epsilon^{\prime}\leq\epsilon$. The specific calculation is as follows:
@@
Random Perturbation Flip. Existing work [45, 46] shows that searching on the nodes of the $l_{\infty}$ ball is more efficient than searching in the $l_{\infty}$ ball. Motivated by these observations, both [33] and [35] suggest some strategies based on searching on the $l_{\infty}$ ball. Here, we choose [33] to further accelerate the attack. After the frequency noise projection, we randomly partial coordinates to flip the signs of perturbation. Suppose $s\in\{-1,1\}^{d}$ and $p\in(0,1)^{d}$, the random perturbation flip is formulated by:
@@
The frequency decision-based attack is shown in Algorithm 3. Based on the cross-task perturbation initialization, the frequency decision-based attack can threaten face forgery detection and better ensure the image quality, which is convenient for the application of downstream tasks.
Algorithm 3 Frequency Decision-based Attack
Input: Fake face $x$, real face $x_{r}$, face recognition model $Z$, face forgery detector $f$
Parameters: Maximum perturbation $\epsilon$, step size $\gamma$
Output: Adversarial face $x_{adv}$
1:  #$\mathrm{\#\ Perturbation\ Initialization}$
2:  $\delta_{init}\leftarrow\mathrm{\textbf{CrossPerturbInit}}(x,x_{r},Z,f)$
3:  $\delta\leftarrow\mathrm{\textbf{BinarySearch}}(x,\delta_{init})$
4:  $\epsilon^{\prime}\leftarrow||\delta||_{\infty}$
5:  while $\epsilon<\epsilon^{\prime}$ do
6:     $x_{dct}^{p}\leftarrow\mathcal{D}(x+\delta)$$\mathrm{\#\ Frequency\ Noise\ Projection}$
7:     $\eta\sim\{-\gamma,\gamma\}^{d}$
8:     $x_{idct}^{p}\leftarrow\mathcal{D’}(x_{dct}^{p}+\eta)$
9:     $x^{p}\leftarrow\mathrm{Clip}_{x,\ \epsilon^{\prime}-\kappa}(x_{dct}^{p})$
10:     if $c(x^{p})=0$ then
11:        $\delta\leftarrow x^{p}-x$
12:     end if
13:     $s\sim\mathrm{\textbf{Bernoulli}}(p)$$\mathrm{\#\ Perturbation\ Random\ Flip}$
14:     $\delta^{p}\leftarrow\delta\odot s$
15:     if $c(x+\delta^{p})==0$ then
16:        $\delta\leftarrow\delta^{p}$
17:     end if
18:     $\epsilon^{\prime}\leftarrow||\delta||_{\infty}$
19:  end while
20:  $x_{adv}\leftarrow x+\delta$
21:  return  $x_{adv}$
IV Experiments
IV-A Experimental Setup
Datasets. FaceForensics++ (FFDF) [18] is a popular large-scale face forgery dataset containing 1,000 video sequences. It includes four forging methods, Face2Face [10], FaceSwap [11], DeepFakes [12] and NeuralTextures [13]. We randomly select 4 frames of each video in the test set, a total of 560 frames (140$\times$4=560). CelebDF [19] is a large-scale DeepFake video dataset with a total of 5,639 DeepFake videos. We also randomly select 500 frames from different videos from the testset.
[…]
TABLE I: The accuracy of spatial-based and frequency-based face forgery detectors on the CelebDF and FFDF datasets.
[…]
@@ IV-A Experimental Setup
Attack Methods. We compare the proposed algorithm with state-of-the-art attack algorithms, including Boundary Attack [29], Opt-Attack [30], Sign-Opt [31], Triangle Attack (TA) [34], HSJA [32], SignFilp [33], and RayS [35]. Attack parameters are consistent with corresponding attacks.
Evaluation metrics. We choose three metrics to quantitatively evaluate the performance of attack methods: attack success rate (ASR), average queries (AQ), and median queries (MQ). Attack success rate (ASR) is defined as the proportion of faces that are successfully attacked. Here, a successful attack is to classify fake faces as real faces. Average queries (AQ) and median queries (MQ) are only calculated over successful attacks, following [33] and [35].
Implementation details. We use Retinaface [48] to detect and crop faces and follow the open-source PyDeepFakeDet [^1] https://github.com/wangjk666/PyDeepFakeDet/blob/main/DATASET.md for preprocessing. For the cropped face, the model resizes it to an image shape and then uses it as the model’s input. For decision-based attacks, we set the maximum perturbation of each pixel to be $\epsilon=0.05$ and the maximum query budget is 10,000. Since these attacks are initialized with randomness, it is easy to cause the attack to fail, so we choose to use the same real face as the initialization if the attack allows. For cross-task perturbation initialization, the cosine similarity of intermediate layer features is interpolated to the same dimension. Here, $K=10$, $\gamma=1.75$, $\xi=0.031$, and $p=0.999$. In the frequency decision-based attack, the attack ends when $\epsilon\geq\epsilon^{\prime}$ or the query reaches the maximum number of queries.
TABLE II: Attack results of decision-based attacks against spatial-based face forgery detection on FaceForensics++ and CelebDF.
@@
TABLE III: Attack results of decision-based attacks against frequency-based detectors on FaceForensics++ and CelebDF.
@@
FFDF Boundary 3313.41 2456.0 12.06 2202.50 1427.0 3.04
Opt 3435.38 3286.0 5.53 3549.99 3029.0 79.46
SignOpt 4262.19 3956.0 19.56 1078.89 785.0 45.90
TA 1000.00 1000.0 33.21 1000.00 1000.0 77.14
HSJA 199.38 187.0 82.50 886.46 188.0 81.42
SignFlip 732.11 280.0 93.57 102.07 86.0 100.00
RayS 140.30 117.0 99.92 225.33 157.0 100.00
Ours 152.25 114.0 100.00 79.19 60.0 100.00
CelebDF Boundary 3393.24 2883.0 22.00 1920.33 14.0 1.80
Opt 1839.61 1178.0 95.60 5053.90 5363.0 30.60
SignOpt 1220.62 868.0 59.20 2752.79 2549.0 71.20
TA 1000.00 1000.0 89.20 1000.00 1000.0 90.00
HSJA 187.04 186.0 100.00 1275.91 324.0 33.60
SignFlip 19.54 12.0 97.20 103.84 12.0 100.00
RayS 84.54 77.0 100.00 84.21 79.0 100.00
Ours 12.59 12.0 100.00 32.67 12.0 100.00
IV-B Attacks on State-of-the-art Detectors

### v2

@@ 3.2. Cross-task Perturbation Initialization
We assume that if there is a high degree of similarity between the intermediate features of the face recognition model and face forgery models, then attacking the face recognition model is equivalent to attacking the face forgery detection models. To verify the correlation of intermediate features across face tasks, we conduct an empirical study on FaceForensics++ [ffpp]. We randomly select 256 frames to calculate the intermediate features and then average for calculating cosine similarity. Figure 2 shows the cosine similarity of intermediate features between face forgery detectors and face recognition models. Here, the face recognition models are FaceNet [FaceNet], CosFace [CosFace], and ArcFace [ArcFace]. The face forgery detection models are ResNet50 [resnet], Xception [xception], EfficientNet-b4 (Eb4) [EfficientNet]. The middle layer selection of face recognition models is shown in Table 7 and the face forgery detectors select the second block unit. According to Figure 2, we find that face recognition models and face forgery detectors have a high correlation between the intermediate features of the same face because these are all face-related tasks and models capture the general features of faces.
@@
Algorithm 1 illustrates the whole pipeline of the cross-task perturbation initialization (CPI), where $\mathrm{sign}$ represents the sign operation on the gradient and $\mathrm{Clip}_{x_{r},\ \xi}(\cdot)$ projects the perturbation $\delta$ to $\xi$-ball in order to satisfy $\delta||_{\infty}\leq\xi$. The time complexity of CPI is $O(K)$, so the calculation of CPI requires $K$ times of forward and backward propagation of the face recognition model. In this paper, $K=10$. We test the time to calculate CPI on a NVIDIA A100, which is about 50.3450.34ms for an image, with almost no additional overhead.
To sum up, we leverage the fast sign gradient to generate perturbations efficiently and focus on the high feature correlation between face-related tasks, which is beneficial to the generalizability of our method to adversarial attacks and defenses on more face-related tasks. As facial recognition models constitute fundamental infrastructure for face-related tasks, they extract the most essential facial features. Face forgery detection, being another face-centric task, is also highly sensitive to these features. If the high correlation between intermediate features of facial recognition and face forgery detection, perturbations designed to attack facial recognition models can consequently disrupt face forgery detection to a significant extent. As convincingly revealed in Figure 2 of the main body, the intermediate features of facial recognition models and face forgery detectors exhibit strong correlation when processing identical face images. Therefore, CPI provides superior initialization for attacking face forgery detection systems. This ensures a more precise optimization trajectory while enabling simultaneous attacks on multiple face-related intelligent systems.
3.3. Frequency Decision-based Attack
Recent work [fre_analysis, f3net, M2TR] on face forgery detection has illustrated a significant difference in frequency between real and fake faces. In addition, some work [highfre, Fourier, SSA] also shows that different models rely on different frequency components to make decisions. Inspired by these works, we consider adding noise directly to the frequency of human faces, aiming to eliminate the difference between real and fake faces in frequency to improve the attack efficiency.
@@ 3.3. Frequency Decision-based Attack
[…]
@@
Based on the general framework of decision-based attacks [hsja, SignFlip], we propose the frequency decision-based attack (FDA) on face forgery detection. Figure 3 summarizes the pipeline of our proposed attack, including the perturbation initialization, the frequency noise projection, and the random perturbation flip. Specifically, we first generate the cross-task adversarial perturbation $\delta_{init}$ by perturbation initialization and resort to binary search to reduce the perturbation magnitude and guarantee adversarial. In frequency noise projection, we add perturbations in the frequency domain and constrain image quality in the spatial domain. Then, we randomly inverse the sign of the partial perturbation, expecting a better adversarial perturbation. Finally, we iterate in frequency noise projection and random perturbation flip until $c(x_{adv})=0$ and $\delta||_{\infty}\leq\epsilon$.
Perturbation Initialization. Generally speaking, the perturbation initialization is used in decision-based attacks to make clean examples keep adversarial. Given a face classified as real, we utilize the cross-task perturbation initialization to obtain initial perturbations $\delta_{init}$, which satisfies the adversarial constraint. Note that cross-task adversarial perturbations are imperceptible, so faces after cross-task perturbation initialization remain adversarial. Further, in order to reduce the magnitude of the initial perturbation, we choose to apply the binary search to make the initial adversarial face as close to $x$ as possible. Binary search in perturbation initialization is summarized in Algorithm 2. Given an initial perturbation $\delta_{init}$, we continuously dichotomize the perturbation range from $0,\max(\delta_{init})]$ on the premise of ensuring adversarial, and finally obtain the final adversarial perturbation $\delta$. In this paper, we set $k$ to 10. With the help of binary search, we obtain the final initial perturbation $\delta$, where $\delta||_{\infty}\leq\epsilon^{\prime}$ and $\epsilon^{\prime}\in(0,\max|\delta_{init}|]$.
Algorithm 3 Frequency Decision-based Attack
Input: Fake face $x$, real face $x_{r}$, face recognition model $Z$, forgery detector $f$
Parameters: Maximum perturbation $\epsilon$, step size $\gamma$
Output: Adversarial face $x_{adv}$
[…]
5:while $\epsilon<\epsilon^{\prime}$ do
6:  $\mathrm{\#\ Frequency\ Noise\ Projection}$
7:  $x_{dct}^{p}\leftarrow\mathcal{D}(x+\delta)$
8:  $\eta\sim\{-\gamma,\gamma\}^{d}$
9:  $x_{idct}^{p}\leftarrow\mathcal{D^{\prime}}(x_{dct}^{p}+\eta)$
10:  $x^{p}\leftarrow\mathrm{Clip}_{x,\ \epsilon^{\prime}-\kappa}(x_{dct}^{p})$
11:  if $c(x^{p})=0$ then
12:   $\delta\leftarrow x^{p}-x$
13:  end if
14:  $\mathrm{\#\ Perturbation\ Random\ Flip}$
15:  $s\sim\mathrm{\textbf{Bernoulli}}(p)$
16:  $\delta^{p}\leftarrow\delta\odot s$
17:  if $c(x+\delta^{p})==0$ then
18:   $\delta\leftarrow\delta^{p}$
19:  end if
[…]
21:end while
[…]
23:return $x_{adv}$
Frequency Noise Projection. In frequency noise projection, we add perturbations in the frequency domain and constrain image quality in the spatial domain. Here, $\mathcal{D}(\cdot)$ denotes discrete cosine transform (DCT) and $\mathcal{D^{\prime}}(\cdot)$ denotes inverse discrete cosine transform (IDCT). First, we transform the image from the spatial domain to the frequency domain with DCT. Then, we add a random noise in the frequency domain, where $\eta\sim\{-\gamma,\gamma\}^{d}$. Finally, we project it to the $\epsilon^{\prime}$-ball in the spatial domain and gradually decrease $\epsilon^{\prime}$ until it is satisfied $\epsilon^{\prime}\leq\epsilon$. To sum up, the specific calculation is as follows:
@@
Random Perturbation Flip. Existing work [PriorConvictions, Parsimonious] shows that searching on the nodes of the $l_{\infty}$ ball is more efficient than searching in the $l_{\infty}$ ball. Motivated by these observations, both [SignFlip] and [RayS] suggest some strategies based on searching on the $l_{\infty}$ ball. Here, we choose [SignFlip] to further accelerate the attack. After the frequency noise projection, we randomly partial coordinates to flip the signs of perturbation. Suppose $s\in\{-1,1\}^{d}$ and $p\in(0,1)^{d}$, the random perturbation flip is formulated by:
@@
The frequency decision-based attack is shown in Algorithm 3 and its time complexity is $O(N)$, where $N$ is the maximum query budget. The main time cost of FDA is the communication and inference time with the API (or offline model). In addition, the calculation of Frequency Noise Projection and Perturbation Random Flip has almost no additional computational cost. Since $N>>K$, the total time complexity is $O(N)$. To sum up, based on the cross-task perturbation initialization, the frequency decision-based attack can threaten face forgery detection and better ensure the image quality, which is convenient for the application of downstream tasks.
Table 1: The accuracy of spatial-based and frequency-based face forgery detectors on the CelebDF and FFDF datasets.
Model ResNet50 Xception Eb4 ViT F3Net M2TR
@@
4. Experiments
4.1. Experimental Setup
[…]
Models. Face forgery detection is based on the spatial domain or frequency domain. For spatial-based face forgery detection, we choose four spatial-based classification networks, such as ResNet50 [resnet], Xception [xception], EfficientNet-b4 (Eb4) [EfficientNet] and ViT-B (ViT) [vit]. For frequency-based face forgery detection, we choose state-of-the-art models, such as F3Net [f3net] and M2TR [M2TR]. All the models are trained according to the corresponding paper settings. The accuracy of these face forgery detectors is shown in Table 1. For face recognition models, we choose FaceNet [FaceNet] to initialize cross-task perturbations. The weights and implementation of FaceNet [^1] https://github.com/ShawnXYang/Face-Robustness-Benchmark/blob/master/RobFR/networks/FaceNet.py are based on Face-Robustness-Benchmark (RobFR) [^2] https://github.com/ShawnXYang/Face-Robustness-Benchmark.
Table 2: Attack results of decision-based attacks against spatial-based face forgery detection on FaceForensics++ and CelebDF.
@@ 4.1. Experimental Setup
Table 3: Attack results of decision-based attacks against frequency-based detectors on FaceForensics++ and CelebDF.
Model F3Net M2TR F3Net M2TR
@@
Boundary 3313.41 2456.0 12.06 2202.50 1427.0 3.04 3393.24 2883.0 22.00 1920.33 14.0 1.80
Opt 3435.38 3286.0 5.53 3549.99 3029.0 79.46 1839.61 1178.0 95.60 5053.90 5363.0 30.60
SignOpt 4262.19 3956.0 19.56 1078.89 785.0 45.90 1220.62 868.0 59.20 2752.79 2549.0 71.20
TA 1143.31 293.0 50.53 837.70 305.0 58.75 179.04 97.0 88.00 220.31 162.0 87.40
HSJA 199.38 187.0 82.50 886.46 188.0 81.42 187.04 186.0 100.00 1275.91 324.0 33.60
SignFlip 732.11 280.0 93.57 102.07 86.0 100.00 19.54 12.0 97.20 103.84 12.0 100.00
RayS 140.30 117.0 99.92 225.33 157.0 100.00 84.54 77.0 100.00 84.21 79.0 100.00
Ours FFDF 152.25 114.0 100.00 79.19 60.0 100.00 CelebDF 12.59 12.0 100.00 32.67 12.0 100.00
Attack Methods. We compare the proposed algorithm with state-of-the-art attack algorithms, including Boundary Attack [boundary], Opt-Attack [OptAttack], Sign-Opt [signopt], Triangle Attack (TA) [TA], HSJA [hsja], SignFilp [SignFlip], and RayS [RayS]. Attack parameters are consistent with corresponding attacks.
Evaluation metrics. We choose three metrics to quantitatively evaluate the performance of attack methods: attack success rate (ASR), average queries (AQ), and median queries (MQ). Attack success rate (ASR) is defined as the proportion of faces that are successfully attacked. Here, a successful attack is to classify fake faces as real faces. Average queries (AQ) and median queries (MQ) are only calculated over successful attacks, following [SignFlip] and [RayS].
Implementation details. We use Retinaface [retinaface] to detect and crop faces and follow the open-source PyDeepFakeDet [^3] https://github.com/wangjk666/PyDeepFakeDet/blob/main/DATASET.md for preprocessing. For the cropped face, the model resizes it to an image shape and then uses it as the model’s input. For decision-based attacks, we set the maximum perturbation of each pixel to be $\epsilon=0.05$ and the maximum query budget is 10,000. Since these attacks are initialized with randomness, it is easy to cause the attack to fail, so we choose to use the same real face as the initialization if the attack allows. For cross-task perturbation initialization, the cosine similarity of intermediate layer features is interpolated to the same dimension. Here, $K=10$, $\gamma=1.75$, $\xi=0.031$, and $p=0.999$. In the frequency decision-based attack, the attack ends when $\epsilon\geq\epsilon^{\prime}$ or the query reaches the maximum number of queries.
Table 4: Attack performance with different maximum perturbation $\epsilon$ on FaceForensics++.
Model Model $\epsilon=0.031$ $\epsilon=0.05$ $\epsilon=0.062$
@@
ResNet50 Boundary 4326.10 3826.0 1.96 6607.10 6363.0 5.50 5224.28 4774.0 5.89
Opt 5848.47 5615.0 22.19 5105.91 4883.0 65.17 4343.85 4026.0 83.93
SignOpt 3065.18 2893.0 6.07 1780.87 1552.0 5.54 1508.39 1133.0 6.07
TA 2050.71 1289.0 57.32 1907.89 981.0 74.11 1829.67 918.0 79.46
HSJA 220.34 190.0 91.07 198.02 190.0 100.00 195.64 190.0 100.00
SignFlip 105.58 98.0 100.00 53.51 50.0 100.00 48.75 40.0 100.00
RayS 213.64 112.0 100.00 152.84 133.0 100.00 141.78 126.0 100.00
Ours 101.38 76.0 100.00 36.16 30.0 100.00 32.29 22.0 100.00
ViT-B Boundary 3328.29 2740.0 5.35 3679.73 3473.0 8.04 3118.15 2890.0 8.39
Opt 5130.87 4622.0 18.51 3393.52 3482.0 30.00 4222.17 3753.0 64.28
SignOpt 2451.80 1872.0 11.17 1342.22 831.0 11.17 1176.37 787.0 11.17
TA 835.56 172.0 72.14 529.64 67.0 91.07 344.77 46.0 95.71
HSJA 288.08 188.0 38.03 245.19 188.0 79.64 264.23 188.0 91.61
SignFlip 644.30 244.0 98.03 167.85 78.0 99.28 105.81 64.0 99.10
RayS 678.24 367.0 100.00 312.35 270.0 100.00 273.83 243.0 100.00
Ours 600.15 330.0 100.00 139.86 42.0 100.00 86.58 42.0 100.00
Table 5: Attack performance of SBI on FaceForensics++.
Model SBI
[…]
Boundary - - 0
Opt - - 0
SignOpt - - 0
TA - - 0
HSJA - - 0
SignFlip - - 0
Rays 544.32 402.0 100.00
Ours 154.23 60.0 100.00
4.2. Attacks on State-of-the-art Detectors